. A turbine, operating under steady-flow conditions, receives 4500 kg of

steam per hour. The steam enters the turbine at a velocity of 2800 m/min, an elevation of 5.5 m

and a specific enthalpy of 2800 kJ/kg. It leaves the turbine at a velocity of 5600 m/min, an elevation

of 1.5 m and a specific enthalpy of 2300 kJ/kg. Heat losses from the turbine to the surroundings

amout to 16000 kJ/h.

Determine the power output of the turbine.



1
Expert's answer
2021-09-13T18:22:37-0400

Gives

Steam rate flow

"\dot{m}=\frac{4500}{3600}=1.25kg/sec"

Steam velocity at inlet

"v_1=\frac{2800}{60}=46.67m/sec"

Velocity of exis "v_2=\frac{5600}{60}=93.34m/sec"

Elevation at inlet "z_1=5.5m"

Elevation at exis "z_2=1.5m"


"Q=16000kJ/hr=\frac{16000}{3600}=4.45KJ/sec"

Enthalpy

"h_1=2800KJ\/kg"

"h_2=2300KJ\/kg"

Turbine operates under steady flow condition


"Q-w_s=h_2-h_1+\frac{1}{2}(v_2^2-v_1^2)+g(z_2-z_1)J/kgsec"

"4.45-w_s=(2300-2800)+\frac{(93.34^2-46.67^2)}{2\times1000}+\frac{9.8(1.5-5.5)}{1000}"

"w_s=4.45-[(2300-2800)+\frac{(93.34^2-46.67^2)}{2\times1000}+\frac{9.8(1.5-5.5)}{1000}]"

"w_s=4.45+500+3.2671-0.0392=507.67KW"